(5,1) is a solution to the equation 2x-3y=7. What does this mean?
Answer by fcabanski(1391)

(5,1) is a coordinate, an ordered pair. Such coordinates/solutions are given in the format (x,y). When x = 5, y=1.


2x-3y=7 --> 2(5) - 3(1) = 10-3 = 7